MOA-II Obsevation Log Monthly Summary 2016/12
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | - | 0 | 0 | 0 |
LMC | 13699-13770 | 72 | 301 | 87400 |
SMC | 1587-1591 | 5 | 10 | 2600 |
TRANSIT | 12587-13138 | 552 | 552 | 16560 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 629 | 863 | 106560 (29h00m2160s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
6 | 39 | 6 | 6 | 6 | 9 | 39 | 39 | 34 | 12 | 9 | 42 |
Field13 | Field14 | Field15 | Field16 |
12 | 42 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
5 | 5 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
TRANSIT FIELD STATISTICS
Field1 | Field2 | Field3 |
107 | 251 | 194 |
SPECIAL TARGET NAME
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2016/12/1: Cloudy most of the night. / night time = 20040(sec)
2016/12/2: We did flat images for Alex's demonstration and the transit survey work / night time = 19860(sec)
2016/12/3: Observations for tr1, tr2, tr3 were carried out in addition to LMC observations / night time = 19680(sec)
2016/12/4: Dome-flat and dark images were taken for transit survey. / night time = 19500(sec)
2016/12/5: Dark images for 300s and 30s were taken. lmc2 was observed but with bad seeing (3.9) and cloudy image for lmc12 / night time = 19380(sec)
2016/12/6: Warning alarm from UPS black box, replacement of its battery is needed / night time = 19200(sec)
2016/12/7: Warning sound from UPS (MGE UPS SYSTEM) in server room / night time = 19080(sec)
2016/12/8: Transit tr2 and tr3, LMC and SMC were observed / night time = 18960(sec)
2016/12/9: CCD temp reading was frozen before 11:30pm because the power supply output was off. The problem might have happened since 07/12 that's why no CCD temp plots of those days were uploaded. / night time = 18780(sec)
2016/12/10: The vacuum value was frozen at 5.9e-06 from 11:30 pm, and I recovered the vacuum reading back to the normal at 12:35 am by switching off and on the vacuum sensor. I also switched off and on the vacuum sensor again at 1:35 am because the values between the sensor and the reading on the computer seemed not matched. / night time = 18720(sec)
2016/12/11: Flat-dome images were taken for transit survey / night time = 18600(sec)
2016/12/12: / night time = 18480(sec)
2016/12/13: / night time = 18420(sec)
2016/12/14: / night time = 18360(sec)
2016/12/15: / night time = 18240(sec)
2016/12/16: / night time = 18180(sec)
2016/12/17: / night time = 18120(sec)
2016/12/18: / night time = 18120(sec)
2016/12/19: / night time = 18120(sec)
2016/12/20: / night time = 18060(sec)
2016/12/21: Dome FLAT, Gain FLAT / night time = 18060(sec)
2016/12/22: / night time = 18060(sec)
2016/12/23: / night time = 18120(sec)
2016/12/24: / night time = 18120(sec)
2016/12/25: / night time = 18120(sec)
2016/12/26: / night time = 18180(sec)
2016/12/27: / night time = 18240(sec)
2016/12/28: Dome FLAT / night time = 18300(sec)
2016/12/29: / night time = 18420(sec)
2016/12/30: / night time = 18480(sec)
2016/12/31: / night time = 18600(sec)
TOTAL NIGHT TIME: 576600(sec)
OPERATING RATE: 24.5 %
GB: 0.0 %
LMC: 17.2 %
SMC: 0.5 %
TRANSIT:6.7 %
ALERT: 0.0 %
GRB: 0.0 %
Return to index
Previous month
Next month